# A Comparative Study of Machine Learning Algorithms for Stock Price Prediction Using Insider Trading Data

### v1

@@ I Introduction
For a long time, financial analysts and investors have been interested in the subject of insider trading [1]. Insider Trading refers to buying or selling securities of a publicly traded company by employees who possess confidential information about that company that has not been made public [2, 3]. Insider trading is regulated by laws, and when insiders deal in equities in accordance with those regulations, their acts are regarded as legal [2]. Insider trading data can deliver useful insights into stock prices. Exploring trends in insider trading data might help determine whether insiders have a favorable or unfavorable outlook on the future of the company [4]. For instance, if a significant number of executives are purchasing their company’s stocks, it can be a sign that they are optimistic about the future performance of the business, which could raise stock prices. On the other hand, if insiders are selling their shares, it can be a sign that they are less optimistic about the future prospects of the business, which could result in a decline in the stock prices of that company. In light of this, examining insider trading data can deliver investors and financial analysts valuable details to help them decide whether to purchase or sell certain stocks.
@@ II Methodology
Figure 1: Bar graphs of Tesla’s monthly average stock prices and monthly total sollar volume traded.

### v2

@@ I Introduction
For a long time, financial analysts and investors have been interested in the subject of insider trading [1]. Insider Trading refers to buying or selling securities of a publicly traded company by employees who possess confidential information about that company that has not been made public [2, 3]. Insider trading is regulated by laws, and when insiders deal in equities in accordance with those regulations, their acts are regarded as legal [2]. Data on insider trading is a major tool for predicting stock prices as it reflects the actions and expectations of those with the most knowledge of the financial condition and future changes of a company [b18].
Exploring trends in insider trading data might help determine whether insiders have a favorable or unfavorable outlook on the future of the company [4]. This type of data offers a distinctive perspective on the mood within the company, which might suggest forthcoming stock price changes before they show in the market [4]. For instance, if a significant number of executives are purchasing their company’s stocks, it can be a sign that they are optimistic about the future performance of the business, which could raise stock prices. On the other hand, if insiders are selling their shares, it can be a sign that they are less optimistic about the future prospects of the business, which could result in a decline in the stock prices of that company. Insider trading may also take place before significant market-moving events like mergers, acquisitions, financial disclosures or leadership changes [b19]. In light of this, analyzing insider trading data can deliver investors and financial analysts valuable details to help them decide whether to purchase or sell certain stocks.
@@ II Methodology
Figure 1: Bar graphs of Tesla’s monthly average stock prices and monthly total dollar volume traded.